OpenStudy (anonymous):

@whpalmer4: That's what I was thinking, but how do you get the measurements?

OpenStudy (anonymous):

You need a number that, when cubed, gives 120. That number would be the cube root of 120. If your calculator doesn't have a cube root button, use 120^(1/3) instead to find 4.93 inches.

OpenStudy (anonymous):

I'll try that. Hang on.

OpenStudy (anonymous):

Hey, it worked. @srossd Thanks.

OpenStudy (anonymous):

Awesome, no problem.

OpenStudy (anonymous):

How would you get a square of one hundred and twenty cubic inches?

OpenStudy (anonymous):

Sorry, I have a daunting homework assignment due after spring break.

OpenStudy (whpalmer4):

Well, you wouldn't get a square with an area of 120 cubic inches. You might get a square with an area of 120 square inches, however. Area of a square with side length \(s\) is\[A=s^2\]You want \(A = 120 \text{ in}^2\)so \[120 \text{ in}^2 = s^2\]Take the square root of both sides: \[\sqrt{120\text { in}^2} = s\]\[s = 2\sqrt{30}\text{ in}\approx 10.9545 \text{ in}\]